# Adaptive Modality Balanced Online Knowledge Distillation for Brain-Eye-Computer based Dim Object Detection

### v1

@@ V-D Adaptive Modality Balancing Analysis
We then analysis the effectiveness of the AMB module and compare the best sensitivity control of the AMB module.
V-D1 The effectiveness of the AMB
@@ V-D2 The sensitivity of the AMB
V-E Generalization Analysis
V-E1 Generalization Analysis on Algorithm Level
As shown in Fig. 13, the unimodality models trained with AMBOKD approach consistently achieve higher performances than the models trained by their own in all combinations. In addition, some uni-modal models, such as the EEGNet model in Fig. 13 (a)$\sim$(b), and the ResNet50 model in Fig. 13 (f), experienced considerable performance improvements compared with their original counterparts after dynamically trained with AMBOKD. For example, EEGNet trained under AMBOKD outperforms the original results in AUC by 7.7 %, ResNet50 trained under AMBOKD shows an improvement of 9.3 % in AUC (see Fig. 13 (a), (c)). These findings indicate the superiority and generalization capabilities of the AMBOKD in multimodal learning tasks.

### v2

@@ V-D Adaptive Modality Balancing Analysis
We then analysis the effectiveness of the AMB module and the influence of hyperparameter $\gamma$ on the AMB module.
V-D1 Effectiveness Analysis
@@ V-D2 Sensitivity Analysis
V-E Generalization Analysis
V-E1 Generalization Analysis on Algorithm Level
@@
As shown in Fig. 13, the unimodality models trained with AMBOKD approach consistently achieve higher performances than the models trained by their own in all combinations. Some of the uni-modal models show significant performance improvements over their original counterparts after training under AMBOKD. For example, EEGNet trained under AMBOKD outperforms the original results in AUC by 7.7 %, ResNet50 trained under AMBOKD shows an improvement in AUC of 9.3 % in AUC (see Fig. 13 (a)$\sim$(c)). These findings indicate the superiority and generalization capabilities of the AMBOKD in multimodal learning tasks.
@@ V-F Physical Verification
Figure 15: The results on physical verification experiment.
VI Conclusion